When purchasing bulk orders of​ batteries, a toy manufacturer uses this acceptance sampling​ plan: Randomly select and test 45 batteries and determine whether each is within specifications. The entire shipment is accepted if at most 2 batteries do not meet specifications. A shipment contains 5000 ​batteries, and 2​% of them do not meet specifications. What is the probability that this whole shipment will be​ accepted? Will almost all such shipments be​ accepted, or will many be​ rejected?
Answer #1

Ans:

Let x be the number of defetive batteries or which do not meet specification.

Then,x has binomial distribution with n=45,p=0.02

P(x=k)=45Ck*0.02k*0.9845-k

P(shipment accepted)=P(x<=2)=P(x=0)+P(x=1)+P(x=2)

=45C0*0.020*0.9845+45C1*0.021*0.9844+45C2*0.022*0.9843

=0.4029+0.3700+0.1661

=0.9390

93.9% shipments will be accepted,6.1% will be rejected.

So,most of the shipment will be accepted and very few will be rejected.